# Magnetocaloric effect measurements in ultrahigh magnetic fields up to 120 T

### v1

@@ 3 Results and discussions
To verify our measurement setup, we first applied a (nondestructive) pulsed magnetic field with a maximum field of T at an initial temperature of K. Figure 2(a) shows the time evolution of the magnetic field and the amplitude of the RF transmission signal obtained by numerical lock-in analysis, respectively. During the field-up sweep ( s), a clear increase in the transmission amplitude is observed. This corresponds to a decrease in the resistance and thus to an increase in the temperature of the thin-film thermometer. Note that the sign of the observed resistance change is opposite to the magnetoresistance (MR) of Au16Ge84 [2]. We expect that the present observation reflects the intrinsic MCE in Ho2Ti2O7, as reported in Ref. [4], where the magnetic entropy is dramatically reduced in association with the magnetic-structure change from the 2-in–2-out state to the 3-in–1-out (1-in–3-out) state. Subsequently, the transmission amplitude decreases and exhibits a dip (indicated by an arrow) slightly after s, where the field reverses its sign. The deviation of the timing of the dip from is 200 ns, which can be attributed to a delay of temperature response.
Figure 2(b) shows the MCE data measured in destructive pulsed fields up to 120 T at three different initial temperatures. Although initial electrical noise results in a poor single-to-noise ratio immediately after the discharge, the transmission signal becomes clear after 1.5 s. At K, the temperature increase during the field-up sweep is more pronounced than that observed in the T shot. For all initial temperatures, the temperature increase tends to saturate above 100 T, consistent with the saturation of the magnetization [4]. During the field-down sweep, in addition to the dip observed slightly after , another dip appears around 50–100 T (indicated by open triangles). This characteristic feature is likely correspond to a decrease in the sample temperature caused by an increase in magnetic entropy associated with the crystal-field level crossing expected in this field range [4].

### v2

@@ 3 Results and discussions
To verify our measurement setup, we first applied a (nondestructive) pulsed magnetic field with a maximum field of T at an initial temperature of K. Figure 2(a) shows the time evolution of the magnetic field and the amplitude of the RF transmission signal obtained by numerical lock-in analysis, respectively. During the field-up sweep ( s), a clear increase in the transmission amplitude is observed. This corresponds to a decrease in the resistance and thus to an increase in the temperature of the thin-film thermometer. Subsequently, the transmission amplitude decreases and exhibits a dip (indicated by an arrow) slightly after s, where the field reverses its sign. The absence of a clear temperature increase during the field down-sweep suggests that the influence of eddy-current heating is not significant. We also note that the sign of the observed resistance change is opposite to the magnetoresistance (MR) of Au16Ge84 [2]. Accordingly, the present observation would reflect the intrinsic MCE in Ho2Ti2O7, as reported in Ref. [4], where the magnetic entropy is dramatically reduced in association with the magnetic-structure change from the 2-in–2-out state to the 3-in–1-out (1-in–3-out) state. The deviation of the timing of the dip from is 200 ns, which can be attributed to a delay of temperature response.
Figure 2(b) shows the MCE data measured in destructive pulsed fields up to 120 T at three different initial temperatures. Although initial electrical noise results in a poor signal-to-noise ratio immediately after the discharge, the transmission signal becomes clear after 1.5 s. At K, the temperature increase during the field-up sweep is more pronounced than that observed in the T shot. For all data sets, the temperature increase tends to saturate above 100 T, consistent with the saturation of the magnetization [4], and exhibits a dip immediately after . The reason why the dip is observed most clearly for K is unclear. In addition to the dip observed slightly after , another dip appears around 50–100 T during the field-down sweep (indicated by open triangles). This characteristic feature is likely correspond to a decrease in the sample temperature caused by an increase in magnetic entropy associated with the crystal-field level crossing expected in this field range [4].